1.57 What is the maximum distance from which a sailor on top of the mast of ship 1, at 34 m above the ocean’s surface, can see another sailor on top of the mast of ship 2, at 26 m above the ocean’s surface?


This problem is too hard.

In the student's manual solutions there is 39km as an answer for this problem.


Please EXPLAIN to me why is the distance 39km.


I think the distance is close to 29km.


and


I got a different answer that the distance is 10km


Please provide me the answer.

THANK YOU VERY MUCH.


1
Expert's answer
2021-11-22T12:38:15-0500

"R=6371000\ (m)"


"l=x_1+x_2"


"x_1=(6371000+34)^2-6371000^2=20814\ (m)"


"x_2=(6371000+26)^2-6371000^2=18201\ (m)"


"l=20814+18201=39015\ (m)\approx39\ (km)" . Answer